Eqn x square + kx + cx=0 has equal roots and one roo of the eqn x sqaure + kx -12=0 is 2 then (k,c)=?

Eqn  x square + kx + cx=0 has equal roots and one roo of the eqn x sqaure + kx -12=0 is 2 then (k,c)=?
 

Grade:12th pass

2 Answers

Imran khan
5 years ago
x^2+kx-12=0 has root:-2 ...................(given)
=>-k+under root k^2+48whole divide2=2
=>by solving this we get k=4....put this in below equation
x^2+kx+cx=0,,,,then we get
x^2 +4x+cx=0....and also given that this equation has same roots then;
we know that;
D=0
=>b^2-4ac=0
4^2-4.1.c=0
16-4c=0
4c=16
=>c=4
therefore values of c and k are 4 and 4....
 
Imran khan
5 years ago
oh sorry here …
D=0
b=4+c
=>4+c whole square -0=0
16+c^2+8c=0
by solving this we get
c=-4  and we had previous got that k=4